OpenStudy (osanseviero):

Before i go sleeping, i knew a formula or logic before of 1 + 1 = 3. It was something like 4 - 10 = 9 - 15 4 -10 + 6 1/4 = 9-15 + 6 1/4 And then more equations...it was complex, and i want to remember how was it to get there to the 2 = 3 ▌1+1 = 3.▐

OpenStudy (anonymous):

well, 1 + 3 is not 3 so you probably used some flawed logic to reach that result. my guess is you wrote 4 - 5 = 9 - 10 (true) then 4 - 10 + 6 + 1/4 = 9 - 15 + 6 + 1/14 then (2 - 5/2)^2 = (3 - 5/2)^2 then you took the square root of both sides: 2 - 5/2 = 3 - 5/2 then you added 5/2 to both sides: 2 = 3 hence 1 + 1 = 3 the flaw of this logic is when you take the square root of both sides...the truth is that the square root of x^2 is the absolute value of x, and it's not necessarily x. so the square root of (2 - 5/2)^2 will be the absolute value of 2 - 5/2, which is in fact 1/2, not -1/2

OpenStudy (anonymous):

the very top line in my last post should have said "1 + 1 is not 3"...

OpenStudy (osanseviero):

thanks. You made my day